OpenStudy (anonymous):

Determine the slope of the line that contains the points E(6,3) & F(-6,3). Select one: a. -1/2 b. 0 c. 1/2 d. undefined

OpenStudy (javalos01):

hey

OpenStudy (anonymous):

hola

OpenStudy (javalos01):

Hablas espaniol

OpenStudy (anonymous):

no

OpenStudy (javalos01):

Yo hablo ingles y espaniol

OpenStudy (javalos01):

ohh

OpenStudy (javalos01):

u dont

OpenStudy (anonymous):

no

OpenStudy (javalos01):

anyways i know the answer

OpenStudy (anonymous):

haha what is it?

OpenStudy (javalos01):

I cant tell direct answers in front of other users

OpenStudy (javalos01):

Ill tell u

OpenStudy (anonymous):

then tell me fam

OpenStudy (javalos01):

Think negative with positive is allways a negative sooo

OpenStudy (javalos01):

A.-1/2

OpenStudy (anonymous):

negative 1/2

OpenStudy (javalos01):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (javalos01):

call me when u need help ok

OpenStudy (anonymous):

alrighty

OpenStudy (javalos01):

Fan me please

OpenStudy (javalos01):

I'll fan u ok

OpenStudy (javalos01):

Hey

OpenStudy (anonymous):

done

OpenStudy (javalos01):

u know how i got 47 sp's

OpenStudy (javalos01):

i was at 20

OpenStudy (anonymous):

good, i just started this soo im just tryna get math answers

OpenStudy (javalos01):

ok im a smart math guy

OpenStudy (javalos01):

what grade r u in

OpenStudy (anonymous):

do you know this one? Choose an equation in slope-intercept form for the line that passes through (0,7) and is parallel to the y=4x-19. Select one: a. y= (1/4)x + 9 b. y = 4x + 7 c. y= (1/4)x - 7 d. y = 4x + 3

OpenStudy (anonymous):

im a junior

OpenStudy (javalos01):

sure give me a minute

OpenStudy (javalos01):

hahaha im so smart

OpenStudy (javalos01):

guess what grade i'm in

OpenStudy (anonymous):

what grade?

OpenStudy (javalos01):

6th

OpenStudy (anonymous):

hahaha oh wow

OpenStudy (javalos01):

im smart huh

OpenStudy (anonymous):

yeah ,i hate math

OpenStudy (javalos01):

i know al lot of things

OpenStudy (javalos01):

dont worry thats what freinds are for

OpenStudy (anonymous):

hahah thank you

OpenStudy (javalos01):

ill help in that question

OpenStudy (anonymous):

okay

OpenStudy (javalos01):

its b

OpenStudy (anonymous):

thank you!!

OpenStudy (javalos01):

your welcome

OpenStudy (javalos01):

i cant believe i know this stuff

OpenStudy (javalos01):

now u want to fan me

OpenStudy (anonymous):

i did!!

OpenStudy (javalos01):

cool i did also

OpenStudy (anonymous):

okay thankss

OpenStudy (javalos01):

im on all day exept suday and sat

OpenStudy (anonymous):

thats good to know. can i ask you another?

OpenStudy (javalos01):

i need to go fast ok

OpenStudy (javalos01):

i'll come back ok

OpenStudy (anonymous):

alright!

OpenStudy (javalos01):

im back

OpenStudy (javalos01):

im gonna help other people alright just mention my name if u need help.ok

OpenStudy (anonymous):

okay!

OpenStudy (javalos01):

im gonna go to other problems